# Multimodal ArXiv: A Dataset for Improving Scientific Comprehension of Large Vision-Language Models

### v2

@@ A.1 Caption Cleaning
Before Cleaning
After Cleaning
@@
A.2 Caption Word Cloud
We visualize the word cloud of captions in our ArXivCap dataset in Figure 7, it can be seen that the captions have a diverse vocabulary for describing the different figures in the academic papers.
A.3 Illustration Cases from ArXivCap and ArXivQA
@@
Figure 10: A case from ArXivQA.
Figure 11: A case from ArXivQA.
Figure 12: A case from ArXivQA.
Figure 13: A case from ArXivQA.
@@ B.1 Details of Evaluated Models
BLIP2 (Li et al., 2023b), introduces a lightweight Q-Former designed to bridge modality gaps and leverages frozen LLMs. Leveraging LLMs, BLIP-2 can conduct zero-shot image-to-text generation using natural language prompts. We select the BLIP2-OPT-6.7B version for evaluation. [^2] https://huggingface.co/Salesforce/blip2-opt-6.7b
@@
InstructBLIP (Dai et al., 2023) employs an instruction-aware visual feature extraction module based on BLIP2 (Li et al., 2023b) and is trained with unified multimodal instruction tuning datasets. We choose InstructBLIP-Vicuna-7B for evaluation. [^3] https://huggingface.co/Salesforce/instructblip-vicuna-7b.
@@
LLaVA (Liu et al., 2023b), adopts Vicuna models as the backbone LLM and is trained on the ChatGPT/GPT-4 generated instruction tuning dataset. LLaVA-v1.5 (Liu et al., 2023a) improves on LLaVA models by employing curated task datasets and an enhanced modality alignment module. We evaluate both LLaVA-v1.5-7B [^4] https://huggingface.co/liuhaotian/llava-v1.5-7b and LLaVA-v1.5-13B. [^5] https://huggingface.co/liuhaotian/llava-v1.5-13b
@@
Flamingo (Alayrac et al., 2022) pioneers the development of LVLMs by introducing a cross-gated layer for LLMs to produce visual-grounded text. The training dataset consists of interleaved visual data and text from the web pages, enabling it to generate free-form text as the output. We select the open-source implementation OpenFlamingo-9B (Awadalla et al., 2023) for evaluation. [^6] https://huggingface.co/openflamingo/OpenFlamingo-9B-vitl-mpt7b
@@
IDEFICS is another open-sourced implementation of Flamingo (Alayrac et al., 2022) Trained on publicly available image-text alignment pairs and instruction tuning datasets, it demonstrates comparable results with the original closed-source model on various image-text benchmarks. We select the IDEFICS-Instruct-9B for evaluation. [^7] https://huggingface.co/HuggingFaceM4/idefics-9b-instruct.
@@
Qwen-VL-Chat (Bai et al., 2023b) is a bilingual LVLM that supports both English and Chinese built on the Qwen LLM (Bai et al., 2023a). During the training phase, Qwen-VL-Chat adopts a packing strategy to create multiple images as inputs, improving its ability to understand the vision context. We select the Qwen-VL-Chat-7B for evaluation. [^8] https://github.com/QwenLM/Qwen-VL
@@
GPT-4V  (OpenAI, 2023), the proprietary vision-language models developed by OpenAI, which are shown to be powerful on various multi-modal tasks (Yang et al., 2023a). The API version we queried is gpt-4-vision-preview.
@@
Bard (Google, 2023) (now Gemini), a commercial LVLM developed by Google. We utilize the unofficial API [^9] https://github.com/dsdanielpark/Bard-API querying the model with our task prompts, accessed on 2023-11-17.
@@ B.2 Task Prompts
  • Multiple-Figure Captioning: Create a caption for the provided figures.
@@ C.1 Caption Type Illustration
We illustrate captions of four types in our main paper in Figure 14. The Acceptable caption provides a comprehensive description of the figure presented. The oversimplified caption is too short compared with the original human-written caption. Furthermore, as show in the third block in Figure 14, Contextual Misinterpretation refers to captions with unmentioned content in the figure, such as the dataset colored in red. Recognition Error denotes the model wrongly identified the number or text in the figure, such as the misidentified model name in the last block of Figure 14.
@@ C.2 Failure Sample of MathVista
Figure 15: A failure case on the geometry problem-solving task.

### v3

@@ A.1 Caption Cleaning
Before Cleaning After Cleaning
@@
A.2 Illustration Cases of Multimodal ArXiv
@@
A.3 Caption Word Cloud
We visualize the word cloud of captions in our ArXivCap dataset in Figure 9. It can be seen that the captions have a diverse vocabulary for describing the different figures in the academic papers.
Figure 10: A case from our ArXivQA dataset.
Figure 11: A case from our ArXivQA dataset.
Figure 12: A case from our ArXivQA dataset.
Figure 13: A case from our ArXivQA dataset.
@@ A.4 ArXivQA Prompting Template
A.5 Quality Analysis of ArXivQA
1. Ensuring Factual Integrity and Clear Presentation: - Factual Alignment: Ensure questions and options are grounded in accurate reflections of the chart data. - Visual Clarity: Maintain high-resolution charts to ensure that all pertinent details are discernible. - Unambiguous Textual Information: Employ precise and unambiguous language to formulate questions and answers, thereby mitigating potential misinterpretations. 2. Ensuring Relevance and Integrated Comprehensiveness: - Question and Option Relevance: Charts must align with their questions, and all options should be applicable and relevant to the given data. - Comprehensive Integration: Guarantee the provision of comprehensive information necessary for the interpretation of the chart and the resolution of the question, ensuring a cohesive amalgamation of textual and visual data. 3. Promoting Fairness and Avoiding Bias: - Equitable Content: Strive for impartiality in the dataset to prevent bias and ensure fair representation of diverse groups and perspectives. Grading Protocol: Each criterion is to be rigorously evaluated for each dataset entry. The assessment is to be conducted on a qualitative scale with three distinct levels: High, Medium, and Low. These levels will denote the degree of conformity to the respective criterion: - 1: High: The dataset entry exhibits exemplary adherence to the evaluation criterion, demonstrating a robust and comprehensive alignment with the specified standard. - 0.5: Medium: The dataset entry meets the evaluation criterion to a moderate extent, indicating a satisfactory but not optimal congruence with the standard. - 0 Low: The dataset entry falls short of the evaluation criterion, signaling a need for significant improvements to meet the standard.
Table 12: Annotator guideline of ArXivQA manual quality examination.
Aspect Avg Score
Factual Alignment 0.6975
Visual Clarity 0.9925
Unambiguous Textual Information 0.9825
Question and Option Relevance 0.9375
Comprehensive Integration 0.905
Equitable Content 1.0
Score Sum 5.515
Table 13: Manual quality analysis of ArXivQA. The average scores for each aspect are presented.
To evaluate the quality of ArXivQA, we manually assess it from six different aspects. We develop a quality examination guideline for annotators, as shown in Table 12, which addresses various aspects of the QA pairs. We sample 100 examples and ask four authors to conduct the quality analysis. The four authors are divided into two groups, with each group tasks with evaluating 50 examples across six sub-aspects, according to the grading protocol.
The evaluation results are presented in Table 13. We find that most samples feature clear, high-quality images with clear and high-quality images, with unambiguous question and option descriptions. However, a small fraction of the generated questions may be unanswerable due to mis-recognizing elements in the figures, as reflected by lower factual alignment scores. Additionally, we consider samples with an aggregate score of 5 or higher from both annotators to be of sufficient quality. Under this stringent criterion, 79 out of 100 samples meet the threshold, demonstrating that the dataset’s quality is generally satisfactory.
@@ B.1 Details of Evaluated Models
(Li et al., 2023b), introduces a lightweight Q-Former designed to bridge modality gaps and leverages frozen LLMs. Leveraging LLMs, BLIP-2 can conduct zero-shot image-to-text generation using natural language prompts. We select the BLIP2-OPT-6.7B version for evaluation. [^3] https://huggingface.co/Salesforce/blip2-opt-6.7b
InstructBLIP
@@
(Dai et al., 2023) employs an instruction-aware visual feature extraction module based on BLIP2 (Li et al., 2023b) and is trained with unified multimodal instruction tuning datasets. We choose InstructBLIP-Vicuna-7B for evaluation. [^4] https://huggingface.co/Salesforce/instructblip-vicuna-7b.
LLaVA
@@
(Liu et al., 2023b), adopts Vicuna models as the backbone LLM and is trained on the ChatGPT/GPT-4 generated instruction tuning dataset. LLaVA-v1.5 (Liu et al., 2023a) improves on LLaVA models by employing curated task datasets and an enhanced modality alignment module. We evaluate both LLaVA-v1.5-7B [^5] https://huggingface.co/liuhaotian/llava-v1.5-7b and LLaVA-v1.5-13B. [^6] https://huggingface.co/liuhaotian/llava-v1.5-13b
Flamingo
@@
(Alayrac et al., 2022) pioneers the development of LVLMs by introducing a cross-gated layer for LLMs to produce visual-grounded text. The training dataset consists of interleaved visual data and text from the web pages, enabling it to generate free-form text as the output. We select the open-source implementation OpenFlamingo-9B (Awadalla et al., 2023) for evaluation. [^7] https://huggingface.co/openflamingo/OpenFlamingo-9B-vitl-mpt7b
Annotation Instruction: As an annotator, your role is to serve as an unbiased and objective judge in evaluating the accuracy of captions produced by a Large Vision-Language Model (LVLM) for scientific figures. These figures are extracted from academic papers, and to aid your assessment, we will provide you with the paper’s title and abstract for necessary context. You will be presented with the original caption—referred to as the ’ground truth’—and the LVLM generated caption, termed the ’prediction’. You could take into account the context given by the paper’s title and abstract for background knowledge, comparing it critically with both captions. In your assessment, please pay attention to the factual alignment, including but not limiting to the following aspects: - Numerical data and statistics: Verify their accuracy and correspondence to the data presented in the figure. - Symbols: Check for correct representation and usage in the context of the scientific subject matter. - Factual content: Ensure all facts are consistent with those stated in the ground truth caption and the paper’s content. <title>title</title> <abstract>abstract</abstract> <ground truth>gt</ground truth> <prediction>pred</prediction> Compare the prediction to the ground truth, provide a brief analysis, and assign a score using one of the following quality labels: <Perfect>, <Good>, <Fair>, <Poor>, <Incorrect>. Below we describe the detail criteria for score: <Perfect>: The prediction is almost identical to the ground truth, with only minor, inconsequential differences that do not change the meaning. All numerical data, symbols, and factual content are accurate and consistent. <Good>: The prediction is largely similar to the ground truth but has some noticeable differences that may slightly change the meaning. However, the core information is still correct, and the numerical data, symbols, and factual content are mostly accurate and consistent with the figure content. <Fair>: The prediction captures the basic idea of the ground truth but has significant differences that change the meaning in a way that cannot be ignored. There may be some inaccuracies or inconsistencies in the numerical data, symbols, or factual content when compared to the figure content. <Poor>: The prediction is related to the ground truth but has serious errors or omissions that significantly change the meaning. The numerical data, symbols, or factual content may be largely inaccurate or inconsistent with the figure content. <Incorrect>: The prediction is completely different or irrelevant to the ground truth, with no similarities between the two. The numerical data, symbols, and factual content are entirely inaccurate or inconsistent with the figure content. Give a brief analysis within 100 words and then output a quality label wrapped with "<>".
Table 14: Prompt template designed for GPT-4 to evaluate generated captions based on the paper title, abstract, and ground truth.
Model BLEU-2 ROUGE-L BERT-S GPT-4 Score
BLIP-2-OPT-6.7B 1.5 6.6 81.3 1.18
InstructBLIP-Vicuna7B 3.5 10.3 83.6 1.48
LLaVA-1.5-7B 2.3 10.4 83.3 1.80
LLaVA-1.5-13B 2.7 11.0 83.6 1.69
OpenFlamingo-9B 5.8 10.3 82.7 1.52
IDEFICS-Instruct-9B 2.1 9.3 83.8 1.55
Qwen-VL-Chat 4.7 11.1 82.0 1.81
Qwen-VL-Chat tuned w/ ArXivCap 8.6 15.3 83.2 2.03
Table 15: Results of 500 single-figure captions generated by various models.
IDEFICS
@@
is another open-sourced implementation of Flamingo (Alayrac et al., 2022). Trained on publicly available image-text alignment pairs and instruction tuning datasets, it demonstrates comparable results with the original closed-source model on various image-text benchmarks. We select the IDEFICS-Instruct-9B for evaluation. [^8] https://huggingface.co/HuggingFaceM4/idefics-9b-instruct.
Qwen-VL-Chat
@@
(Bai et al., 2023b) is a bilingual LVLM that supports both English and Chinese built on the Qwen LLM (Bai et al., 2023a). During the training phase, Qwen-VL-Chat adopts a packing strategy to create multiple images as inputs, improving its ability to understand the vision context. We select the Qwen-VL-Chat-7B for evaluation. [^9] https://github.com/QwenLM/Qwen-VL
GPT-4V
@@
(OpenAI, 2023), the proprietary vision-language models developed by OpenAI, which are shown to be powerful on various multi-modal tasks (Yang et al., 2023a). The API version we queried is gpt-4-vision-preview.
Bard
@@
(Google, 2023), a commercial LVLM developed by Google. We utilize the unofficial API [^10] https://github.com/dsdanielpark/Bard-API querying the model with our task prompts, accessed on 2023-11-17.
Gemini 1.0 Pro Vision
(Reid et al., 2024), a upgraded LVLM by Google. We utilize the official API querying the model with our task prompts, accessed on 2024-05-20.
@@ B.2 Task Prompts
Multiple-Figure Captioning Create a caption for the provided figures.
@@
B.3 GPT-4 Evaluation of Caption
In addition to BLEU-2, ROUGE-L, and BERT-S, we also utilize GPT-4 to evaluate a sample of 500 generated captions. Specifically, we employ GPT-4 for the evaluation of single-figure caption tasks following FairEval (Wang et al., 2023). The template for prompting GPT-4 to evaluate generated captions is presented in Table 14. GPT-4 is asked to perform an analysis and then produces a quality score, which is subsequently mapped to a scale from 1 to 5. The results are presented in Table 15. We observe that the ROUGE-L metric exhibits the highest correlation with the GPT-4 Score (Pearson r = 0.91), followed by BLEU-2 (Pearson r = 0.64). BERT-S instead demonstrates a moderate correlation (Pearson r = 0.39). The uniformly low GPT-4 scores across all models suggest that they struggle to produce satisfactory captions, which is consistent with the findings in our main paper. Notably, training on ArXivCap results in a significant 12% improvement in the GPT-4 score compared to the original Qwen-VL-Chat model, leading to the most favorable outcomes in this evaluation.
@@ Appendix C Error Analysis
Figure 15: A failure case on the geometry problem-solving task.
@@ C.1 Caption Type Illustration
We illustrate captions of four types in our main paper in Figure 14. The Acceptable caption provides a comprehensive description of the figure presented. The oversimplified caption is too short compared with the original human-written caption. Furthermore, as shown in the third block in Figure 14, Contextual Misinterpretation refers to captions with unmentioned content in the figure, such as the dataset colored in red. Recognition Error denotes the model wrongly identified the number or text in the figure, such as the misidentified model name in the last block of Figure 14.
@@ C.2 Failure Sample of MathVista
Appendix D Results with LLaVA Backbone
Model MathVista MMMU(val) ScienceQA(IMG only) MM-Vet
LLaVA-v1.5-7B 26.6 35.3 66.8 30.5
   Original SFT +ArXivQA 28.2 36.0 68.3 32.4
Table 16: After fine-tuning with a combination of ArXivQA and original SFT data, the LLaVA model shows boosted mathematical reasoning abilities across benchmarks.
We investigate whether ArXivQA could also enhance other LVLMs, such as LLaVA models (Liu et al., 2023b). To maintain model performance, we mix our ArXivQA dataset with the LLaVA SFT 665K-instruction tuning dataset. The LLaVA-v1.5-7B is adopted as the backbone and the model is trained following the original recipe. The results on various benchmarks are listed in Table 16. We find that not only the scientific reasoning performance is improved on multimodal reasoning tasks (MathVista (Lu et al., 2023), MMMU (Yue et al., 2023), and ScienceQA (Lu et al., 2022)), but the overall capability on MM-Vet (Yu et al., 2024) is also boosted. Together with our results using Qwen-VL-Chat, these findings indicate that our ArXivQA dataset can enhance different model backbones and is beneficial across various benchmarks.